OpenStudy (anonymous):

Suppose a rocket ship in deep space moves with constant acceleration equal to 9.8 m/s2, which gives the illusion of normal gravity during the flight. (a) If it starts from rest, how long will it take to acquire a speed one-tenth that of light, which travels at 3.0 × 108 m/s? (b) How far will it travel in so doing?

OpenStudy (anonymous):

errr your way off

OpenStudy (anonymous):

lol much simpler than that

OpenStudy (anonymous):

a is in metres and b is in seconds for the answers

OpenStudy (anonymous):

sorry you have misinterpreted the question

sam (.sam.):

whats the answer?

OpenStudy (anonymous):

a) 10.6m (b)41.5s

OpenStudy (anonymous):

i have no idea about the working lol

sam (.sam.):

The acceleration is constant, and =g, in the moving (accelerating) frame, which means that dα/dτ = g/c, where β = v/c = tanh α α = the velocity parameter τ = proper time g = acceleration of gravity at Earth's surface =~9.80 m/s^2 v = velocity c = speed of light in vacuum =~3.00*10^8 m/s tanh is the hyperbolic tangent, defined in terms of hyperbolic sine and cosine: tanh x = sinh x / cosh x; where sinh x = (e^x - e^-x)/2 cosh x = (e^x + e^-x)/2, so that tanh x = (e^x - e^-x)/(e^x + e^-x) = (e^2x - 1)/(e^2x + 1) = 1 - 2/(e^2x + 1) (a) This actually has two answers -- elapsed time aboard ship (τ), and elapsed time back on Earth (t) So, because at t=0, τ, v, β, and α are all =0, dα/dτ = g/c means that α = gτ/c For v = 0.12c, β = 0.12, α = arctanh β = 0.12058, so τ = cα/g = 3.691*10^6 s = 42.72 days =======================PROPER TIME==================== =6 weeks + 17 hr, as experienced in the spaceship (cause it's proper time, which goes with the moving object) To get elapsed time, t, in the "Earth" frame, use the definition of proper time in terms of "Earth" spacetime coordinates, t and x: dτ^2 = dt^2 - (dx/c)^2; dτ = √(1 - β^2)dt = dt/γ where γ = 1/√(1 - β^2) = cosh α dα/dτ = g/c then becomes g/c = γdα/dt = cosh α dα/dt the solution of which is sinh α = gt/c; ---> from which we get: t = (c/g)sinh α = γv/g = v/(g√(1 - β^2)) = 3.700*10^6 s = 42.83 days ===============COORDINATE TIME ON EARTH=================== =~6 weeks + 20 hr, as seen on Earth (b) To get distance traveled (this has to be in the "Earth" frame): α = arcsinh(gt/c) β = tanh α = sinh α/cosh α = sinh α/√[1 + (sinh α)^2] = gt/√[c^2 + (gt)^2] v = dx/dt = cβ = cgt/√[c^2 + (gt)^2], which we integrate, and use the fact that x=0 when t=0 (that's where the "-1" comes from at the end of the next line): x = (c^2/g)(√[1 + (gt/c)^2] - 1) = 6.6845*10^13

OpenStudy (anonymous):

OHHH

OpenStudy (anonymous):

sorry i was looking at the wrong question window...ignore my last replies...however..

OpenStudy (anonymous):

(a) 3.1 × 106 s; (b) 4.6 × 1013 m

OpenStudy (anonymous):

10^6 not 106

OpenStudy (anonymous):

10^13

sam (.sam.):

τ = cα/g = 3.691*10^6 s

sam (.sam.):

For distance i get x= 6.6845*10^13m

OpenStudy (anonymous):

hmm ok this is just a physics problem (motion in 1d)

OpenStudy (anonymous):

it can be done just using constant acceleration formulas....don't need trig

OpenStudy (anonymous):

1st year uni stuff...maybe high school stuff

sam (.sam.):

Without the relativistic (SR) realm, using newtonian kinematics, this question can be done too, vf = 0.12 * 3 * 10^8 m/sec Given v0 = 0 and a = 9.8 m/sec^2 At constant acceleration, vf = v0 + at vf = 9.8 m/sec^2 t Equate the two speeds vf = 9.8t = 0.12 ( 3 ) x10^8 m/sec. t = 3.673469.3x10^6 secs. for (b) From the time computed above, compute the distance at constant acceleration and v0 = 0 using the distance formula: s = 1/2 at^2 s = 6,612x10^13 m.

OpenStudy (anonymous):

ah k answers still different

OpenStudy (anonymous):

i'm getting it from a textbook :p

OpenStudy (anonymous):

thanks for trying